OpenStudy (anonymous):

Find the critical numbers of the function. h(t) = t^3/4 − 3t^1/4

OpenStudy (anonymous):

What did you get for the derivative?

OpenStudy (anonymous):

I no im supposed to find the derivative

OpenStudy (anonymous):

and solve for 0, but i didnt no how to solve for 0 for this equation

OpenStudy (anonymous):

i got 3/4t^-1/4-3/4t^-3/4

OpenStudy (anonymous):

Looks right. Now set that = 0.

OpenStudy (anonymous):

i did that but because of the exponenti get confused

OpenStudy (anonymous):

\[\frac{3}{4t^{1/4}} - \frac{3}{4t^{3/4}} = 0\]

OpenStudy (anonymous):

so i factor out 3/4t^-1/4?

OpenStudy (anonymous):

I wouln't bother. Just set \(t^{1/4} = t^{3/4}\)

OpenStudy (anonymous):

Cause that's the only way those two things can be equal.

OpenStudy (anonymous):

\[\frac{3}{4}t^{-1/4} - \frac{3}{4}t^{-3/4} = 0\] \[\frac{3}{4}(t^{-1/4} - t^{-3/4}) = 0\] \[t^{-1/4} - t^{-3/4} = 0\] \[t^{-1/4} = t^{-3/4}\] \[t^{1} = t^{3}\] \[\implies t=1\]

OpenStudy (anonymous):

how did u get t^1=t^3

OpenStudy (anonymous):

raised both sides to the -4 power.

OpenStudy (anonymous):

ooooo so it get crossed out

OpenStudy (anonymous):

omg this way is much simpler then the way my teacher explained it!

OpenStudy (anonymous):

thank u soo much!

OpenStudy (anonymous):

But really you coulda guessed t would be 1 looking at the original equation for the derivative because \[\frac{3}{4t^a} - \frac{3}{4t^{3a}} = 0\] Will only be true if \(t^a = t^{3a}\) which means that t must be 1.

OpenStudy (anonymous):

o ya i c that now